2. A uniform charge of Q is placed on the sphere of a Van de Graff generator, such that electron discharge is observed from a conducting point a distance of 46 cm away (center to center). Let the dielectric strength of air... E-3 x 10 N/C Determine the Electric Potentials at the 50 cm and 80 cm radial locations a) b) A small air balloon of charge q8.4 nC and mass 1.76 grams, is placed at rest at the 80 cm potential surface, and then released. Determine its speed v at the 50 cm radial location: (neglect any air resistance) Hint: Work-Energy Theorem… W = AK Scanned with CamScanner
Answer #1

2

a)

Q = charge = ?

r = distance = 0.46 m

E = dielectric strength = 3 x 106 N/C

Using the equation

E = k Q/r2

3 x 106 = (9 x 109) Q/(0.46)2

Q = 70.5 x 10-6 C

a)

r1 = 50 cm = 0.50 m

electric potential is given as

V1 = k Q/r1 = (9 x 109) (70.5 x 10-6)/(0.50) = 1.3 x 106 Volts

r2 = 80 cm = 0.80 m

electric potential is given as

V2 = k Q/r2 = (9 x 109) (70.5 x 10-6)/(0.80) = 7.93 x 105 Volts

b)

m = mass = 1.76 g = 1.76 x 10-3 kg

q = charge = - 8.4 x 10-9 C

using conservation of energy

qV2 = q V1 + (0.5) m v2

(- 8.4 x 10-9) (7.93 x 105) = (- 8.4 x 10-9) (1.3 x 106 ) + (0.5) (1.76 x 10-3 ) v2

v = 2.2 m/s